A sample of xenon gas at a pressure of 954 mm Hg and a temperature of 72°C, occupies a volume of 7.20 liters. If the gas is heated at constant pressure to a temperature of 104 °C, the volume of the gas sample will be?
Answer #1

Charles' law states that at constant pressure, the volume occupied by a gas is directly proportional to its absolute temperature.

Mathematically,

\frac{V}{T} = Constant

Or, for a gas at 2 different conditions of volume and temperature,

\frac{V1}{T1} = \frac{V2}{T2}

In the given question, let V1 = 7.20 L,

T1 = 72 ℃ = 345 K,

T2 = 104 ℃ = 377 K.

So, \frac{7.20 L}{345 K} = \frac{V2}{377 K}

=> V2 = \frac{7.20 L \times 377K }{345 K}

=> V2 = 7.87 L.

Therefore, the required volume is 7.87 L.